OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (6, 1) (5, 4) is it y= 2x+ 5 ?

OpenStudy (anonymous):

m = (4-1)/(5-6) = 3/(-1) = -3 y - 1 = -3 (x - 6) y = -3 (x -6) +1 y = -3x + 18 +1 y = -3x + 19